Rules of the Game

Srinivas Palanki

In most matters it is more important that the applicable rule of law be settled than that it be settled right

Course Instructor

Louis D. Brandeis Dr. Srinivas Palanki Email: palanki@eng.fsu.edu Website: www.eng.fsu.edu/ palanki Email is the easiest way of contacting the instructor Dr. Srinivas Palanki Email: palanki@eng.fsu.edu Website: www.eng.fsu.edu/ palanki Email is the easiest way of contacting the instructor Dr. Srinivas Palanki Email: palanki@eng.fsu.edu Website: www.eng.fsu.edu/ palanki Email is the easiest way of contacting the instructor Dr. Srinivas Palanki Email: palanki@eng.fsu.edu Website: www.eng.fsu.edu/ palanki Email is the easiest way of contacting the instructor TEXT BOOK

1. Analysis, Synthesis and Design of Chemical Processes, 2nd Ed., R. Turton, R. C. Bailie, W. B. Whiting, and J. A. Shaeiwitz, Prentice Hall PTR, Upper Saddle River, New Jersey 07458, 2003.

\underline{SOFTWARE}

1. CHEMCAD

2. Excel

Please contact Dr. Palanki to get a copy of CHEMCAD.

\underline{REFERENCE BOOKS}: Textbooks used in (1) Mass & Energy Balances (2) Thermodynamics, (3) Separations Processes, (4) Kinetics & Reactor Design.

Rules of the Game – p.4/9

TEXT BOOK

1. Analysis, Synthesis and Design of Chemical Processes, 2nd Ed., R. Turton, R. C. Bailie, W. B. Whiting, and J. A. Shaeiwitz, Prentice Hall PTR, Upper Saddle River, New Jersey 07458, 2003.

\underline{SOFTWARE}

1. CHEMCAD

2. Excel

Please contact Dr. Palanki to get a copy of CHEMCAD.

\underline{REFERENCE BOOKS}: Textbooks used in (1) Mass & Energy Balances (2) Thermodynamics, (3) Separations Processes, (4) Kinetics & Reactor Design.

Rules of the Game – p.4/9

TEXT BOOK

1. Analysis, Synthesis and Design of Chemical Processes, 2nd Ed., R. Turton, R. C. Bailie, W. B. Whiting, and J. A. Shaeiwitz, Prentice Hall PTR, Upper Saddle River, New Jersey 07458, 2003.

\underline{SOFTWARE}

1. CHEMCAD

2. Excel

Please contact Dr. Palanki to get a copy of CHEMCAD.

\underline{REFERENCE BOOKS}: Textbooks used in (1) Mass & Energy Balances (2) Thermodynamics, (3) Separations Processes, (4) Kinetics & Reactor Design.

Rules of the Game – p.4/9

TEXT BOOK

1. Analysis, Synthesis and Design of Chemical Processes, 2nd Ed., R. Turton, R. C. Bailie, W. B. Whiting, and J. A. Shaeiwitz, Prentice Hall PTR, Upper Saddle River, New Jersey 07458, 2003.

\underline{SOFTWARE}

1. CHEMCAD

2. Excel

Please contact Dr. Palanki to get a copy of CHEMCAD.

\underline{REFERENCE BOOKS}: Textbooks used in (1) Mass & Energy Balances (2) Thermodynamics, (3) Separations Processes, (4) Kinetics & Reactor Design.

Rules of the Game – p.4/9

Prerequisites for ECH 4604

ECH 3274L Measurements & Transport Phenomena Lab.
ECH 4267 Advanced Transport Phenomena
ECO 2023 Principles of Microeconomics
Corequisite ECH 4504 Kinetics & Reactor Design

Students not meeting these prerequisites must drop the course. Courses for which ECH 4604 is a Prerequisite

ECH 4323 Process Control
ECH 4323L Process Control Laboratory
ECH 4615 Chemical Engineering Process Design II

The above courses are taught only in the spring semester. Students failing ECH 4604 will likely set back their progress towards graduation by one year.

Prerequisites for ECH 4604

ECH 3274L Measurements & Transport Phenomena Lab.
ECH 4267 Advanced Transport Phenomena
ECO 2023 Principles of Microeconomics
Corequisite ECH 4504 Kinetics & Reactor Design

Students not meeting these prerequisites must drop the course. Courses for which ECH 4604 is a Prerequisite

ECH 4323 Process Control
ECH 4323L Process Control Laboratory
ECH 4615 Chemical Engineering Process Design II

The above courses are taught only in the spring semester. Students failing ECH 4604 will likely set back their progress towards graduation by one year.

Prerequisites for ECH 4604

ECH 3274L Measurements & Transport Phenomena Lab.
ECH 4267 Advanced Transport Phenomena
ECO 2023 Principles of Microeconomics
Corequisite ECH 4504 Kinetics & Reactor Design

Students not meeting these prerequisites must drop the course. Courses for which ECH 4604 is a Prerequisite

ECH 4323 Process Control
ECH 4323L Process Control Laboratory
ECH 4615 Chemical Engineering Process Design II

The above courses are taught only in the spring semester. Students failing ECH 4604 will likely set back their progress towards graduation by one year.

Prerequisites for ECH 4604

ECH 3274L Measurements & Transport Phenomena Lab.
ECH 4267 Advanced Transport Phenomena
ECO 2023 Principles of Microeconomics
Corequisite ECH 4504 Kinetics & Reactor Design

Students not meeting these prerequisites must drop the course. Courses for which ECH 4604 is a Prerequisite

ECH 4323 Process Control
ECH 4323L Process Control Laboratory
ECH 4615 Chemical Engineering Process Design II

The above courses are taught only in the spring semester. Students failing ECH 4604 will likely set back their progress towards graduation by one year.

Course Focus

Course Learning Objectives

Part I: Development of Process Flowsheets Part II: Analysis of Process Flowsheets Part III: Equipment Sizing and Costing Part IV: Analysis of Non-traditional Processes Part I: Development of Process Flowsheets Part II: Analysis of Process Flowsheets Part III: Equipment Sizing and Costing Part IV: Analysis of Non-traditional Processes Part I: Development of Process Flowsheets Part II: Analysis of Process Flowsheets Part III: Equipment Sizing and Costing Part IV: Analysis of Non-traditional Processes Part I: Development of Process Flowsheets Part II: Analysis of Process Flowsheets Part III: Equipment Sizing and Costing Part IV: Analysis of Non-traditional Processes Students will apply engineering principles to design chemical processes.

Students will analyze the feasibility of unit operations based on thermodynamics, transport, and reaction engineering principles.

Students will synthesize a chemical process from its component parts.

Students will apply economic and engineering principles for estimating equipment costs, capital costs and manufacturing costs.

Students will utilize modern simulation software.

Students will be able to say that “this is the best engineering course that I have taken”.

Rules of the Game – p.7/9

Students will apply engineering principles to design chemical processes.

Students will analyze the feasibility of unit operations based on thermodynamics, transport, and reaction engineering principles.

Students will synthesize a chemical process from its component parts.

Students will apply economic and engineering principles for estimating equipment costs, capital costs and manufacturing costs.

Students will utilize modern simulation software.

Students will be able to say that “this is the best engineering course that I have taken”.

Rules of the Game – p.7/9

Students will apply engineering principles to design chemical processes.

Students will analyze the feasibility of unit operations based on thermodynamics, transport, and reaction engineering principles.

Students will synthesize a chemical process from its component parts.

Students will apply economic and engineering principles for estimating equipment costs, capital costs and manufacturing costs.

Students will utilize modern simulation software.

Students will be able to say that “this is the best engineering course that I have taken”.

Rules of the Game – p.7/9

Students will apply engineering principles to design chemical processes.

Students will analyze the feasibility of unit operations based on thermodynamics, transport, and reaction engineering principles.

Students will synthesize a chemical process from its component parts.

Students will apply economic and engineering principles for estimating equipment costs, capital costs and manufacturing costs.

Students will utilize modern simulation software.

Students will be able to say that “this is the best engineering course that I have taken”.

Rules of the Game – p.7/9

Students will apply engineering principles to design chemical processes.

Students will analyze the feasibility of unit operations based on thermodynamics, transport, and reaction engineering principles.

Students will synthesize a chemical process from its component parts.

Students will apply economic and engineering principles for estimating equipment costs, capital costs and manufacturing costs.

Students will utilize modern simulation software.

Students will be able to say that “this is the best engineering course that I have taken”.

Rules of the Game – p.7/9

Students will apply engineering principles to design chemical processes.

Students will analyze the feasibility of unit operations based on thermodynamics, transport, and reaction engineering principles.

Students will synthesize a chemical process from its component parts.

Students will apply economic and engineering principles for estimating equipment costs, capital costs and manufacturing costs.

Students will utilize modern simulation software.

Students will be able to say that “this is the best engineering course that I have taken”.

Rules of the Game – p.7/9

All course material will be available on BLACKBOARD. Student Teams:

Students will work on teams of 4 for some (but not all) assignments.

Teams will be made by the second day of class.

Team members should sit together in class.

Please review the honor code in the student handbook Grading Weight:

Design Project  40 points
Midterm  25 points
Quizzes & Assignments 35 points

All course material will be available on BLACKBOARD. Student Teams:

Students will work on teams of 4 for some (but not all) assignments.

Teams will be made by the second day of class.

Team members should sit together in class.

Please review the honor code in the student handbook Grading Weight:

Design Project  40 points
Midterm  25 points
Quizzes & Assignments 35 points

All course material will be available on BLACKBOARD. Student Teams:

Students will work on teams of 4 for some (but not all) assignments.

Teams will be made by the second day of class.

Team members should sit together in class.

Please review the honor code in the student handbook Grading Weight:

Design Project  40 points
Midterm  25 points
Quizzes & Assignments 35 points

All course material will be available on BLACKBOARD. Student Teams:

Students will work on teams of 4 for some (but not all) assignments.

Teams will be made by the second day of class.

Team members should sit together in class.

Please review the honor code in the student handbook Grading Weight:

Design Project  40 points
Midterm  25 points
Quizzes & Assignments 35 points

Grading Scale:

85-100  A
70-84  B
50-69  C
40-49  D
less than 40 F

This fixed grading scale will be used. There will be no “grading on the curve”.

Midterm Exam and Design Project Schedule

October 11, 2004  Midterm
November 15, 2004 Project due

Grading Scale:

85-100  A
70-84  B
50-69  C
40-49  D
less than 40 F

This fixed grading scale will be used. There will be no “grading on the curve”.

Midterm Exam and Design Project Schedule

October 11, 2004  Midterm
November 15, 2004 Project due

Grading Scale:

85-100  A
70-84  B
50-69  C
40-49  D
less than 40 F

This fixed grading scale will be used. There will be no “grading on the curve”.

Midterm Exam and Design Project Schedule

October 11, 2004  Midterm
November 15, 2004 Project due

Additional Policies

Attendance policy Bonus points policy Downloading lectures Computer use during class Ethics, professionalism and honor code violations Office hours Attendance policy Bonus points policy Downloading lectures Computer use during class Ethics, professionalism and honor code violations Office hours Attendance policy Bonus points policy Downloading lectures Computer use during class Ethics, professionalism and honor code violations Office hours Attendance policy Bonus points policy Downloading lectures Computer use during class Ethics, professionalism and honor code violations Office hours Attendance policy Bonus points policy Downloading lectures Computer use during class Ethics, professionalism and honor code violations Office hours Attendance policy Bonus points policy Downloading lectures Computer use during class Ethics, professionalism and honor code violations Office hours